FORMATION OF CH^{+}: SHOCK CHEMISTRY IN NGC 7027

\underline{JULIE K. ANDERSON}, Department of Chemistry and Biochemistry, The University of Arizona, Tucson, AZ 85719; LUCY M. ZIURYS, Department of Chemistry and Biochemistry, The University of Arizona, Tucson, AZ 85719; FABRICE HERPIN, LAB-OASU, France.

The formation of CH^{+} in the interstellar medium has been an enigma for the past 70 years. Emission from the species is found in diffuse material, even though the major pathway leading to the species, C^{+} + H_{2}, is endothermic by 0.40 eV. The barrier for this reaction can be greatly reduced if the H_{2} is vibrationally excited. Using the Heterodyne Instrument for the Far Infrared (HIFI) on board the Herschel Space Observatory, we have mapped CH^{+} (J = 1 → 0), CH^{+} (J = 2 → 1), and C^{+} (J = 3/2 → 1/2) across the face of the young planetary nebula, NGC 7027. Analysis of the spectra has shown that CH^{+}, C^{+}, and vibrationally excited H_{2} apparently trace the same outflow in NGC 7027. Therefore CH^{+} in this nebula likely forms from the activated C^{+} + H_{2} reaction. Spectral maps, temperatures, and abundances of CH^{+} will be presented along with spectral maps of C^{+}.